A plane flies with a velocity of 100km/hr to n a direction N60°E. Find its velocity in the easterly and northerly directions



1
Expert's answer
2020-09-06T17:16:29-0400


Here the length of velocity vector is "|\mathbf{v}| = 100\, km/h".

A projection of the velocity "\mathbf{v}" in the easterly direction is


"v_E = |\mathbf{v}|\cos30\degree = 100\cdot \dfrac{\sqrt{2}}{2}\approx 70.71\, km/h"

A projection of the velocity "\mathbf{v}" in the northerly direction is:


"v_N = |\mathbf{v}|\sin30\degree = 100\cdot \dfrac{1}{2}= 50\, km/h"

Answer. "v_E = 70.71\, km/h", "v_N = 50\, km/h".